OpenStudy (anonymous):

find the volume of the region bounded by the graphs y = sqrt(x), y = 0, and x = 6, revolved about the line x = 6

OpenStudy (amistre64):

:) again :)

OpenStudy (anonymous):

hello! This problem is similar to the one we did together yesterday, but for some reason I am still doing this wrong

OpenStudy (anonymous):

I know that I am supposed to do the disk method, and that I need to solve for each equation in terms of y

OpenStudy (amistre64):

rewrite y= sqrt(x) in terms of x = f(y) y = sqrt(x) ; ^2 both side y^2 = x this we can solve for easily right?

OpenStudy (anonymous):

so I have v = pi (the integral from 0 to 6) (y+9)^2 - (9)^2 dx

OpenStudy (amistre64):

graph this and see that it is the same

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

x = 6 is a pain, we want to move it to the x=0 line to examine it; how do we move x = 6 to x=0? x= 6 ; -6 from both sides x-6 = 0 do the same for the x=y^2 to keep it in the same position

OpenStudy (anonymous):

minus 6 from both sides in the equation I wrote?

OpenStudy (amistre64):

x = y^2 ; -6 both sides x-6 = y^2 -6

OpenStudy (anonymous):

whoops---I am revolving around x = 9, sorry about that

OpenStudy (amistre64):

brb..

OpenStudy (anonymous):

okay

OpenStudy (amistre64):

back...

OpenStudy (anonymous):

hi

OpenStudy (amistre64):

since it 9 we just -9 fro itall

OpenStudy (anonymous):

so v = pi (intergral from 0 to 6) (y + 9)^2 - 9^2 ?

OpenStudy (amistre64):

move it all to the left by 9 to examine it at x=0; does that make sense? Its over on a shelf and we want to examine it at the table where we can see it clearly; the table is the origin of the graph where x=0 and y=0, so we move it

OpenStudy (anonymous):

So now I am looking at a graph that is shifted backwards by 9 on my calculator

OpenStudy (amistre64):

y = sqrt(x) ; change to an equal form to accomodate for the y axis y^2 = x ; now move it over to the x=0 by subtracting 9 y^2 -9 = x-9 ; the x-9 can be rewritten as just x

OpenStudy (anonymous):

sqrt(x +9)

OpenStudy (anonymous):

okay so it is y^2 -9 not +9

OpenStudy (amistre64):

sqrt(x+9) is good; now reform it y = sqrt(x+9) y^2 = x+9 y^2 -9 = x

OpenStudy (amistre64):

whats our bounds now: for y=0 to how far?

OpenStudy (anonymous):

x = 6....?

OpenStudy (anonymous):

I've never seen anyone do transformations of a graph this way..

OpenStudy (amistre64):

y = sqrt(9) y = 3

OpenStudy (anonymous):

we have to redo them when we solve for a different variable...keep forgetting

OpenStudy (amistre64):

integrate the discs from 0 to 3

OpenStudy (anonymous):

okay

OpenStudy (amistre64):

pi {S} y^2-9 dy ; [0,3] = volume of revolution

OpenStudy (amistre64):

forgot to square the function lol

OpenStudy (amistre64):

pi {S} [y^2-9]^2 dy ; [0,3]

OpenStudy (anonymous):

so you don't have to subtract another 9?

OpenStudy (amistre64):

why would you subtract two 9s?

OpenStudy (amistre64):

we only move the graph over to x=0 from x=9 right?

OpenStudy (anonymous):

one from the first eqn and one from eqn y = 0?

OpenStudy (amistre64):

if i tellyou to move 9 steps to the left; what do you do? move 18?

OpenStudy (anonymous):

ohhhh

OpenStudy (amistre64):

y=0 is already bounding the volume; no need to adjust that part lol

OpenStudy (anonymous):

answer is 648/5?

OpenStudy (amistre64):

dunno; lets check..

OpenStudy (anonymous):

no, its' not

OpenStudy (amistre64):

pi {S} (y^2-9)^2 dy pi {S} (y^4 -18y^2 +81) dy pi(y^5/5 -18y^3/3 +81y) at y=3

OpenStudy (anonymous):

had that part

OpenStudy (anonymous):

and got 648/5, but the homework grader marked it wrong

OpenStudy (amistre64):

i get 129.6pi

OpenStudy (anonymous):

same answer, and the wrong one. What did we do wrong?

OpenStudy (anonymous):

Hang on. You're rotating the graph of y= sqrt(x) from x=6 to x=9 about the line x=9 ? Is that right?

OpenStudy (amistre64):

243 27 243 pi( ---- - --- + ---) right? 5 3 1

OpenStudy (amistre64):

x = 0 to 9

OpenStudy (anonymous):

right, and we are rotating the region about the line x = 9

OpenStudy (amistre64):

ilove; retype the problem in its entirety so that we have a clear picture

OpenStudy (anonymous):

Find the volumes of the solids generated by revolving the regions bounded by the graphs of the equations about the given lines. y = sqrt(x) y = 0 x = 6 about the line x = 9

OpenStudy (amistre64):

ohhh... so its a donut shape; i assumed you meant y=0; x=9 about the x=9

OpenStudy (anonymous):

oh, nope sorry about that

OpenStudy (anonymous):

I don't think it's a doughnut

OpenStudy (amistre64):

its a torus alright :)

OpenStudy (anonymous):

More like a mountain.

OpenStudy (amistre64):

we have to subtract the middle of what we found to get the volume then

OpenStudy (anonymous):

oh wait. I'm looking at wrong graph

OpenStudy (anonymous):

Small hill. Not a doughnut. There is no hole in the middle.

OpenStudy (amistre64):

x=6 to x=9 is empty tho

OpenStudy (anonymous):

there is no hole it the middle? isn't the shape a bunt cake?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

graph y = sqrt(x) from x = 6 to x=9 y = 0 to y = 3

OpenStudy (anonymous):

but the region is cut off at x = 6, and we have to get it to revolve about line x = 9

OpenStudy (anonymous):

that's where it starts

OpenStudy (anonymous):

not where it ends. it doesn't say it's bounded by x=0

OpenStudy (anonymous):

hmmm...then how to we caluclate where it begins and ends?

OpenStudy (anonymous):

If the problem said bounded by x=0, x=6 and y= 0 and y = sqrt(x) I would agree with that picture. It doesn't say x=0.

OpenStudy (anonymous):

so using what you just said, the boundries would be?

OpenStudy (anonymous):

x = 6, x = 9, y = 0, y = sqrt(x) and rotate around x = 9

OpenStudy (amistre64):

then we would double our results right?

OpenStudy (anonymous):

no, you have the wrong part of the graph I think.

OpenStudy (amistre64):

with this boundary; y = [0,sqrt(6)]

OpenStudy (anonymous):

You want to integrate over the white part in your picture rather than the grey part.

OpenStudy (anonymous):

what wouuld the equation look like?

OpenStudy (anonymous):

\[v = \pi \int\limits_{0}^{3}...\]

OpenStudy (anonymous):

You're using disks or shells?

OpenStudy (anonymous):

We can only use the disk method, right?

OpenStudy (anonymous):

where did you get the boundaries 0 to 3 from , again?

OpenStudy (anonymous):

do you have the value for the answer so we can be sure we're doing the right region?

OpenStudy (anonymous):

nope----online h/w problem

OpenStudy (anonymous):

dumb.

OpenStudy (anonymous):

I'm pretty sure I'm reading it right though. For this I would use cylindrical shells

OpenStudy (amistre64):

y = sqrt(x); curve of sqrt(x) y = 0 ; the x axis is y=0 x = 6 ; the left boundary x=9; axis to spin around

OpenStudy (amistre64):

x=6 the right bound

OpenStudy (anonymous):

x = 6 is the left bound in my interpretation of the problem, and the axis of rotation would be the right bound.

OpenStudy (anonymous):

???? what do you get as your answer?

OpenStudy (anonymous):

I thought the entire region was left bound, and the axis is right bound.

OpenStudy (anonymous):

like left-right or top-bottom

OpenStudy (amistre64):

the region to cut from my original; if im right is: is the cylindar of r = 3, and height = sqrt(6) 216 sqrt(6) pi ------------ 5

OpenStudy (anonymous):

okay, this is the right answer. Can you go over it one more time?